# Rational Metareasoning for Large Language     Models

### v1

@@ 6.2 Reasoning in LLMs
Techniques such as Chain of Thought (CoT) and related methodologies (Wei et al., 2023; Yao et al., 2023; Yasunaga et al., 2024; Madaan et al., 2023; Zheng et al., 2024) have proven effective at enhancing LLM performance across a wide range of tasks. CoT boosts LLMs’ performance on complex reasoning by guiding them through a series of intermediate reasoning steps, increasing inference costs to improve task performance. This method can be implemented through in-context learning (Wei et al., 2023), prompting (Kojima et al., 2023), or training (Li et al., 2023). The benefits of CoT can be attributed to both a greater computation depth (Goyal et al., 2024; Pfau et al., 2024) and the semantic values of the thought tokens, which function as intermediate variables in the computation of the answer (Prystawski et al., 2023). However, recent studies have raised concerns regarding the meaningfulness of such reasoning chains in reaching the target solution, and whether models effectively utilize them to solve tasks (Turpin et al., 2023; Paul et al., 2024; Sprague et al., 2024). We further demonstrate that standard prompting and training methods fail to teach the model to use CoT purposefully, resulting in inefficient inference. Reasoning can also be used to bootstrap language models. Self-improving techniques (Huang et al., 2022; Zelikman et al., 2022) consist of generating reasoning chain-augmented answers for unlabeled questions and fine-tuning the LLM using those self-generated solutions as target outputs. Similar techniques, such as ReST (Gulcehre et al., 2023), can also be used to better align LLMs with human preferences and needs. Our approach builds on these techniques to optimize the inference cost of reasoning chains in addition to their task performance. More recently, chat models that “think before answering” have been developed (OpenAI, 2024), using inference-time computation to enhance their outputs. Although these models outperform others, they expend more computational resources on all tasks, even when it may not be necessary. Our method could be incorporated into their training process to help the model determine when this additional computation is genuinely beneficial.
@@ 7 Conclusion
However, we note several limitations to our approach. First, the focus of our work here is enhancing the efficiency of large language models in reasoning tasks, rather than the overall performance of such systems. Thus, we aimed to reduce unnecessary intermediate reasoning steps, rather than improve the quality of reasoning per se. It remains to be seen whether our approach can be extended to enhance task performance in addition to computational efficiency. Additionally, we tested our approach on well-established datasets in science, commonsense reasoning, and math, but it remains untested in more specialized domains. Broader testing could help assess its generalizability and effectiveness across diverse contexts.
@@ Appendix A Full results tables
A.1 Qualitative examples
Below we show one example of question, reasoning chain and answer per dataset, generated with each method: CoT Few Shots (Yellow), STaR (green) and Metareasoning (Red). The following examples were generated by Phi-2 (Javaheripi et al., 2023).

### v2

@@ 6.2 Reasoning in LLMs
Techniques such as Chain of Thought (CoT) and related methodologies (Wei et al., 2023; Yao et al., 2023a; Yasunaga et al., 2024; Madaan et al., 2023; Zheng et al., 2024) have proven effective at enhancing LLM performance across a wide range of tasks. CoT boosts LLMs’ performance on complex reasoning by guiding them through a series of intermediate reasoning steps, increasing inference costs to improve task performance. This method can be implemented through in-context learning (Wei et al., 2023), prompting (Kojima et al., 2023), or training (Li et al., 2023). The benefits of CoT can be attributed to both a greater computation depth (Goyal et al., 2024; Pfau et al., 2024) and the semantic values of the thought tokens, which function as intermediate variables in the computation of the answer (Prystawski et al., 2023). However, recent studies have raised concerns regarding the meaningfulness of such reasoning chains in reaching the target solution, and whether models effectively utilize them to solve tasks (Turpin et al., 2023; Paul et al., 2024; Sprague et al., 2024). We further demonstrate that standard prompting and training methods fail to teach the model to use CoT purposefully, resulting in inefficient inference. Reasoning can also be used to bootstrap language models. Self-improving techniques (Huang et al., 2022; Zelikman et al., 2022) consist of generating reasoning chain-augmented answers for unlabeled questions and fine-tuning the LLM using those self-generated solutions as target outputs. Similar techniques, such as ReST (Gulcehre et al., 2023), can also be used to better align LLMs with human preferences and needs. Our approach builds on these techniques to optimize the inference cost of reasoning chains in addition to their task performance. Another related work, Quiet-STaR (Zelikman et al., 2024), has improved LLM reasoning performance by modifying pretraining to generate intermediate s between tokens. While effective for downstream tasks, it increases computational costs by generating reasoning chains at every step, even when unnecessary. More recently, chat models that “think before answering” have been developed (OpenAI, 2024), using inference-time computation to enhance their outputs. Although these models outperform others, they expend more computational resources on all tasks, even when it may not be necessary. Our method could be incorporated into their training process to help the model determine when this additional computation is genuinely beneficial.
@@ 7 Conclusion
However, we note some limitations to our approach. First, the focus of our work here is enhancing the efficiency of large language models in reasoning tasks, rather than the overall performance of such systems. Thus, we aimed to reduce unnecessary intermediate reasoning steps, rather than improve the quality of reasoning per se. It remains to be seen whether our approach can be extended to enhance task performance in addition to computational efficiency. Additionally, we tested our approach on well-established datasets in science, commonsense reasoning, and math, but it remains untested in more specialized domains. Notably, one of these untested settings is the agentic one, where LLMs act as agents to perform complex tasks in digital environments (Yao et al., 2023b; Schick et al., 2023). Our method could be adapted by including the number of LLM API calls to such tools in the cost section of our reward function, ideally encouraging the model to minimize unnecessary tool calls. Broader testing could help assess its generalizability and effectiveness across diverse contexts. Additionally, the VoC reward function is highly adaptable and can be used with various algorithms. For instance, it can provide numerical rewards for PPO or guide DPO by training on samples with the highest and lowest rewards. In this work, we focused on EI, the most sample-efficient algorithm for reasoning tasks Havrilla et al. (2024), to demonstrate its generalization across models and benchmarks. Exploring its application to other algorithms is an exciting direction for future research.
@@ Appendix A Full results tables
Appendix B Alternative plot of accuracy vs cost
Appendix C Additional out of domain testing
To further evaluate the out-of-domain generalization capabilities of our method, we conducted additional assessments on AQUA-RAT (Ling et al., 2017) and MMLU Hendrycks et al. (2021) using the Phi-2 model. For MMLU, we incorporated a few-shot prompting baseline with examples drawn from datasets used during training and compared the results to STaR and Metareasoning. AQUA-RAT was tested under the same conditions. The results indicate that our method performs comparably to few-shot prompting and STaR on both datasets, while requiring fewer output tokens. This analysis highlights the strength of our approach in achieving a balance between computational efficiency and accuracy in OOD scenarios. Comprehensive results for MMLU and AQUA-RAT are provided in Table 6.
Dataset Method Accuracy (%) $\uparrow$ # Input Tokens # Output Tokens
MMLU Direct Few-Shot 34.1 ($\pm$ 1.1) 495.7 ($\pm$ 2.1) 0.0 ($\pm$ 0.0)
CoT Few-Shot 50.7 ($\pm$ 1.3) 1050.7 ($\pm$ 2.2) 99.1 ($\pm$ 2.2)
STaR 48.2 ($\pm$ 1.3) 94.7 ($\pm$ 2.1) 110.5 ($\pm$ 1.5)
Metareasoning 48.3 ($\pm$ 1.4) 94.7 ($\pm$ 2.2) 78.7 ($\pm$ 1.6)
Aqua-Rat Direct Few-Shot 22.4 ($\pm$ 5.1) 477.9 ($\pm$ 2.5) 0.0 ($\pm$ 0.0)
CoT Few-Shot 38.6 ($\pm$ 5.5) 1032.9 ($\pm$ 2.6) 152.6 ($\pm$ 11.3)
STaR 42.9 ($\pm$ 5.9) 76.9 ($\pm$ 2.5) 138.8 ($\pm$ 11.5)
Metareasoning 36.6 ($\pm$ 5.9) 76.9 ($\pm$ 2.4) 111.2 ($\pm$ 7.7)
Table 6: Additional OOD testing on Phi-2. This table compares the performance of our method against STaR, CoT and Direct Few-Shot across two out-of-domain benchmarks: MMLU and AQUA-RAT. The results highlight that our method achieves comparable accuracy while requiring fewer output tokens, demonstrating both efficiency and robustness in diverse reasoning tasks.
Appendix D Comparison to instruction tuned models
It is reasonable to question whether instruction-tuned models can dynamically adjust the length of their reasoning chains based on task complexity. To explore this, we tested Meta-Llama-3-8B-Instruct using the following prompt:
Answer the following question, thinking step by step to get to the answer. You can think however long you need, but answer as soon as you’re ready. Keep you response concise and use the minimum number of steps to get to the answer. Once you’re finished thinking, write your answer after the ’Answer: ’ prompt.
Unfortunately, this method did not effectively adapt response lengths to align with task complexity in our experiments, as shown in Tables 7 and 8 and Fig. 4. While the observed performance drop might partially result from deviations from the expected input format, it is evident that the length of the reasoning chains was generally much higher and not meaningfully adaptive to the complexity of the task.
Model Method Accuracy (%) $\uparrow$ # Input Tokens # Output Tokens
Meta-Llama-3-8B Direct Few-Shot 55.5 ($\pm$ 1.2) 421.3 ($\pm$ 1.0) 1.0 ($\pm$ 0.0)
CoT Few-Shot 63.7 ($\pm$ 1.4) 995.0 ($\pm$ 0.9) 148.4 ($\pm$ 3.0)
CoT Instruct 57.7 ($\pm$ 1.3) 73.3 ($\pm$ 0.9) 190.4 ($\pm$ 2.5)
STaR 63.9 ($\pm$ 1.5) 73.3 ($\pm$ 1.0) 119.7 ($\pm$ 1.9)
Metareasoning 64.4 ($\pm$ 1.4) 73.3 ($\pm$ 0.9) 95.6 ($\pm$ 2.3)
Table 7: Llama Instruct results. The table shows accuracy (in %), average number of input tokens, and average number of output tokens for each method. Despite the instruction, reasoning chains remain lengthy.
Figure 4: Llama Instruct plot. Left panel: the distribution of input and output tokens across various methods and datasets. Right panel: Comparison of performance across datasets.
Model Method Hard Split Easy Split Length Reduction (%) $\uparrow$
Meta-Llama-3-8B CoT Few-Shot 132.0 ($\pm$ 3.0) 57.0 ($\pm$ 1.0) 56.8
CoT Instruct 181.0 ($\pm$ 4.0) 150.0 ($\pm$ 2.0) 17.1
STaR 126.0 ($\pm$ 3.0) 64.0 ($\pm$ 2.0) 49.2
Metareasoning 99.5 ($\pm$ 2.5) 27.0 ($\pm$ 1.0) 72.9
Appendix E Qualitative examples
Below we show one example of question, reasoning chain and answer per dataset, generated with each method: CoT Few Shots (Yellow), STaR (green) and Metareasoning (Red). The following examples were generated by Phi-2 (Javaheripi et al., 2023). The reasoning chains shown here serve as intermediate computations that guide the model to its final answers. While not intended as direct explanations, the examples remain interpretable and demonstrate the effectiveness of the respective methods.
@@ Appendix E Qualitative examples
Appendix F Direct Finetuning Baseline
We also tested our method against a direct finetuning baseline. It can be seen that while this method provides an even bigger efficiency gain, the performance is less consistent.
Figure 9: